OpenStudy (anonymous):

a company makes 3 types of cable, CABLE A requires 3 black, 3 white, and 2 red wires. B requires 1 black, 2 white, and 1 red. C requires 2 black, 1 white, and 2 re. They used 100 black, 110 white and 90red wires. How many of each cable were made? a) 10 cable A: 3O cable B; 20 cable C. b) 10 cable ; 103 cable B; 20 cable C. c) 20 cable A; 30 cable B; 10 cable C. d)10 cable A; 30 cable B; 93 cable C. i need the process pleaseee

OpenStudy (anonymous):

yeah yeah hahaha thanks! but i dont know how to replace b can you help mee doing that?

OpenStudy (anonymous):

so the answer is? a, b, c or d?

OpenStudy (anonymous):

Ok, I did that problem totally wrong here u go Let A = number of cable A B = number of Cable B C = number of Cable C # of Black wires used = 3A + 1B + 2C = 100 # of White wires used = 3A + 2B + 1C = 110 # of Red Wires used = 2A + 1B + 2C= 90 you have a system on three equations with three unknowns. Solve the system of equations. solution: A = 10, B = 30, C = 20!

OpenStudy (anonymous):

thats the point idont know how to solve the system of linear equation hahaha:( can you help meee?